OpenStudy (mckenzieandjesus):

Can you please check my answer? Describe how to translate the graph of y= sqrt x to obtain the graph of y=sqrt x-9 shift down 9 units shift left 9 units **** shift right 9 units shift up 9 units

OpenStudy (mckenzieandjesus):

@jim_thompson5910

OpenStudy (defiance):

Whats your answer?

OpenStudy (mckenzieandjesus):

shift left 9 units

OpenStudy (defiance):

Oh. Yes. It's shift left 9 unites. 

OpenStudy (defiance):

It means its moving across the "x" axis. The X is horizontal.

OpenStudy (defiance):

So, you're correct :D.

OpenStudy (mckenzieandjesus):

Okay thanks :)

OpenStudy (jdoe0001):

\(\textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards \\ \quad \\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ % template start y = \sqrt{{\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} )} + {\color{green}{ D}}\\ % template ends \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to\ the\ left \\ \quad \\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to\ the\ right\end{array} \end{array}\)